A ball moves subject to a force that varies with position. The force is FF) = – k = 115 N m², and the ball moves in a straighPlease show calculations :)
Answer #1

w=*** SES ) (36) Ydy 11 9.2 ada [yda W = -k -k - (gign2 + y2) 3/2 G², 2156 ² + 0 2)/2 7 9.2 W = 0 - (U5 N/m2) | 14² +9.52 . W